Sally bought three chocolate bars and a pack of gum and paid $1.75. Jake bought two chocolate bars and four packs of gum and paid $2.00. Find the cost of a chocolate bar and the cost of a pack of gum.
Answer by ptaylor(2198)

let x=cost of chocolate bars
and y=cost of gum
Now we are told that 3x+y=$1.75-------------------------eq1
We are also told that 2x+4y=$2.00-------------------------eq2
Multiply each term in eq1 by 4 and we get:
12x+4y=$7-----------------------------------------new eq1
Subtract eq2 from new eq1 and we have:
10x=$5 divide both sides by 10
x=$0.50----------------------------------cost of chocolate bars
substitute x=$0.50 into eq 2:
2($0.50)+4y=$2 get rid of parens
$1+4y=$2 subtract $1 from both sides
4y=$2-$1
4y=$1 divide both sides by 4
y=$0.25-----------------------------------------cost of gum
CK
3($0.50)+$0.25=$1.75
$1.75=$1.75
and
2($0.50)+4($0.25)=$2
$2=$2
Hope this helps-------ptaylor